OpenStudy (anonymous):

What is the sum of a 6-term geometric series if the first term is 20 and the last term is 336,140?

ganeshie8 (ganeshie8):

try finding the common ratio r, then use series formula

OpenStudy (anonymous):

how do you find the common ratio?

ganeshie8 (ganeshie8):

\(a = 20\) \(t_n = ar^{n-1}\)

ganeshie8 (ganeshie8):

6th term = 336,140 \( t_6 = ar^{6-1}\) \( 336140 = 20r^{5}\)

ganeshie8 (ganeshie8):

you can find r ?

OpenStudy (anonymous):

To find r, do we divide that big number by 20?

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

\(\frac{336140}{20} = r^{5}\)

ganeshie8 (ganeshie8):

you may use your calculator

OpenStudy (anonymous):

so it would be 20 1-16807^6/1-16807?

ganeshie8 (ganeshie8):

close ! but the r you got is incorrect !

ganeshie8 (ganeshie8):

\(\frac{336140}{20} = r^{5}\)\(\) \(16807 = r^5\) \(7^5 = r^5\) \(7 = r\)

ganeshie8 (ganeshie8):

plugin r = 7, in the series formula and find the sum...

OpenStudy (anonymous):

How did you get 7^5 from 16807?

ganeshie8 (ganeshie8):

16807 = 7x7x7x7x7 = 7^5 ive used my calculator

OpenStudy (anonymous):

thank you!

ganeshie8 (ganeshie8):

np :)